OpenStudy (anonymous):

If f(x) = (x+2)/(2), then what is f–1(x)

OpenStudy (anonymous):

f(x) = x^2 /(x-2) We know that if f= u/v , then f'= (u'v-uv')/v^2 Now we will assume that: u= x^2 ==> u'=2x v= x-2 ==> v' =1 Then f'(x) = [2x(x-2)-...

OpenStudy (anonymous):

y = x + 2 / 2 isolate x